OpenStudy (anonymous):

how many solutions will it have write the solution out y=3x-4 x+3y=6

OpenStudy (anonymous):

y=3x-4 write slope intercept -3x+y=-4 time (3) 3x+9y=18 -3x+y=-4 Add------------------- 10y = 14 y=14\10 y= 7\5 subtitutio y=7/5 solve for y x+3*(7/5)=6 x+21/5 = 6 x =6-(21/5) x=(30-21)\5 x=9\5 If I deserve medal then give me one.

OpenStudy (anonymous):

T Y

OpenStudy (anonymous):

yeah